A wallet contains a Php 20-bill, a Php 50-bill, and a Php 100-bill. A bill is taken out from the wallet and then replaced with a bill of the same amount that was taken away. If a bill is taken out from the wallet, what is the probability that the sum of the bill is more than Php 40?


1
Expert's answer
2022-03-02T15:43:49-0500

we define the following

 P(> 40)= P(70) + P(120) + P(150) ="2\times ({1 \over 3 }\times {1 \over 3})+2\times({1 \over 3 }\times {1 \over 3})+2\times({1 \over 3 }\times {1 \over 3})= \dfrac{2}{3}"